# Investigating Solid-Fluid Phase Coexistence in DC Plasma Bilayer Crystals: The Role of Particle Pairing and Mode Coupling

### v1

@@ Abstract
This article presents a detailed investigation of solid-fluid phase coexistence in a bilayer dusty plasma crystal subjected to varying confinement ring bias voltages in a DC glow discharge argon plasma. Melamine formaldehyde particles were employed to form a stable, hexagonally ordered bilayer crystal within a confinement ring electrically isolated from the grounded cathode. By systematically adjusting the confinement ring bias, a distinct phase coexistence emerged: it is characterized by a fluid-like melted core surrounded by a solid crystalline periphery. Crucially, analysis of the phonon spectra revealed frequency shifts that deviate significantly from the predictions of classical monolayer Mode-Coupling Instability (MCI) theory. Stability analysis further demonstrated that dynamic interlayer particle pairing and non-reciprocal interactions play a pivotal role in destabilizing the bilayer structure. These findings highlight previously underappreciated mechanisms driving the melting transition in bilayer dusty plasmas, offering a more comprehensive understanding of phase behavior in complex plasma systems. The results underscore the importance of interlayer coupling and confinement effects in tuning structural transitions.
@@ I Introduction
In recent years, dusty plasma crystals have emerged as a model system for exploring phase coexistence [15, 16, 17, 18, 19, 20, 21], where solid and fluid phases can coexist within the same structure. The dynamics underlying this phenomenon are complex, as they involve particle interactions, structural rearrangements, and instabilities that are not commonly observed in simpler systems. Phonon modes play a critical role in understanding phase transitions and melting dynamics in complex plasmas. Longitudinal and transverse phonon modes, arising from compressional and shear motions respectively, provide insights into system stability and energy distribution [22, 23, 24]. Wang et al. [22] developed a theoretical framework for phonon modes in Yukawa crystals a model analogous to dusty plasmas with screened Coulomb interactions. This enabled systematic studies of dispersion relations and the influence of confinement and interaction strength on phonon behavior and instabilities  [25, 26, 7, 27].
@@
Studies on phase transitions in monolayer complex plasmas have shown that mode-coupling instability (MCI) is the primary mechanism responsible for melting.  [35, 7, 26, 34, 27]. It was observed that the coupling between longitudinal and out-of-plane modes triggers the destabilization of the lattice, ultimately leading to melting. Ivlev et al. [36] later proposed a common framework for monolayer and bilayer crystals, highlighting wake-mediated non-reciprocity, MCI and a universal stability principle for both monolayer and bilayer complex plasmas. The instability is driven by the coupling of wave modes, which is caused by nonreciprocal, wake-mediated interparticle forces. In monolayers, ions wakes introduce a weak negative coupling between the normally separate horizontal and vertical modes. Conversely, bilayers possess an inherent positive coupling between horizontal modes due to their structural asymmetry. Therefore, for a bilayer to become unstable, the wake-induced negative coupling must be strong enough to overcome this existing positive offset, a condition that is facilitated in both systems when modes are in resonance. Recently, Jaiswal et al. [16] reported the first experimental observation of phase coexistence in bilayer dust clouds, induced by confinement variations. While they observed signs of MCI, the absence of intersection between vertical and horizontal modes and phase coexistence seemed to emerge when interlayer and interparticle spacing approach or fall below the screening length. This highlights the need for detailed investigation of wake-mediated non-reciprocal interactions and their role in bilayer crystal stability.
This article investigates the mechanisms underlying phase coexistence and melting in bilayer dusty plasma crystals, with a specific focus on the interplay between particle pairing and mode coupling under varying confinement ring bias voltages in a DC glow discharge plasma. By analyzing phonon spectra across a range of confinement potentials, we explore how longitudinal and transverse modes respond to particle interactions and coupling effects. Through detailed tracking of particle trajectories, separations, and velocities, we uncover new insights into the dynamic processes that destabilize the crystal structure. Our results demonstrate that dynamic particle pairing, combined with a modified form of MCI, serves as a primary driver of phase transitions, ultimately leading to melting. Additionally, we introduce and apply a quantitative analysis of interlayer non-reciprocity, capturing the emergence of asymmetric particle interactions that signal structural instability. This approach provides a novel diagnostic for identifying the onset of disorder and melting in bilayer dusty plasmas. Overall, this study contributes to a deeper understanding of structural stability and phase transitions in complex plasma systems, offering new perspectives on the microscopic interactions that govern these phenomena.
@@ III Results
Fig. 2 shows an overlapped image of 20 consecutive frames from the top view camera at 103 V of confining ring voltage. The circular marked region shows a clear melting whereas the periphery is almost stationary with a small thermal fluctuation around the equilibrium position. The side view of the particle cloud (shown in Fig. 3) reveal that the interlayer gap between two layer is decreasing as the confinement potential decreases from 140 V to 103 V. This increases the interaction between the layers and vertical oscillations and coupling of two layer were seen for the voltage below 120 V (see supplementary material).
@@ III.1 Phonon spectra
The phonon spectrum of the dust lattice [35, 33] was analyzed over a range of confinement ring bias voltages, focusing on optical modes (from side-view data) and acoustic modes (from top-view data). Particle positions were extracted using ImageJ [37] for both views. The extracted positions were processed using the Trackpy library [38, 39] to link particle trajectories and compute velocities. Longitudinal and transverse phonon currents were then calculated following the method of Nunomura et al. [23], and Fourier transforms were applied to obtain the current fluctuation spectra. This analysis across various ring bias voltages enabled us to study the evolution of wave modes in the crystal.
@@
At a confinement bias of 120 V, signs of torsional motion [40] become visible, indicating the onset of particle pairing [see hotspots in Figs. 5(c), (g) and 5(c), (g)]. As previously reported, torsion and MCI develop over similar spatial and temporal scales [40]. However, in contrast to monolayer MCI, no localized hot spots are present in the out-of-plane spectra (Fig. 6(c)). Below 120 V, spectral features appear increasingly blurred. It is possible that the imaging setup captured particles from both layers, leading to artificial mixing of spectral features. This behavior likely reflects a change in interparticle coupling, possibly due to the formation of particle pairs between the two layers. Such pairing can modify the system’s vibrational response and affecting the mechanical stability of the lattice [40].
@@ III.2 Dependence of particle pairing on the confining ring bias voltage
As the ring bias voltage was reduced, a noticeable increase in the number of particle pairs formed was observed. Tracking the number of pairs as a function of voltage revealed a clear trend: the number of pairs increased as the voltage decreased (Fig. 7), particularly around 120 V and below. At 103 V, the number of pairs was measured as 107 ± 10, confirming the strong increase in pair formation at lower confinement potentials
@@ III.3 Case study: dynamics of particle pairing at a confining ring bias voltage of 120 V
As we have seen in previous sections, At 120 V, significant changes in particle behavior were observed. For example, the hot spot arises at 120 V and the pairing effect between the particles also increases drastically after 120 V. This has important consequence on particle structural dynamics therefore 120 V has been considered as a critical voltage for understanding the transition from crystalline to fluid states. In order to illustrate that, we focus on three particles: "1" and "2" in the top layer (Layer 1 in Fig. 8(a)), and "3" just above Layer 1. Their motion over 6 seconds reveals the dynamics of pair formation, collapse, and particle dragging, contributing to system destabilization. Fig. 8(b) shows the evolution of their $x$-coordinates.
@@
Initially, particle "3" formed a stable pair with "1" for $\sim$1.5 s, as shown by their nearly identical $x$-coordinates. At $t\simeq 1.5$ s, the pair broke, and "3" drifted leftward toward particle "2," forming a new pair shortly before $t=2$ s. Around $t\simeq 3$ s, particle "3" continued leftward and while dragging along particle "2" which gradually moved leftwards too. The pair realigned by $t\simeq 4$ s. This dragging is the signature of the ion wake below particle "3" influences nearby particles (also evidence by Melzer et al.[30]), potentially destabilizing the entire system.
@@
where $\gamma$ is the Epstein drag coefficient [41], $\langle V\rangle$ is the average horizontal velocity and $m$ is the mass of the particles. Assuming the system is overdamped and using $F_{x}=k\cdot\Delta x$, the spring constant for the 2-3 pair during the interaction was:
@@
The estimated horizontal dragging force and effective spring constant quantify the strength of the wake-mediated coupling between paired particles. These values confirm that the interparticle attraction is sufficiently strong to produce measurable lateral displacement and energy exchange. The spring constant thus characterizes how mechanical energy is stored and released during pairing events, linking microscopic force asymmetry to the observed velocity fluctuations and, ultimately, to the onset of melting. [34].
@@
III.4 Nonreciprocal interactions and driving strength
@@
To quantify this effect, we introduced a pair-resolved metric following from the recent work from Ivlev et al. [42]
$\mathbf{R}=<|F_{top}+F_{bottom}|>$ (1)
@@
which directly probes the imbalance of forces in paired particles across the two layers. In the case of reciprocal interactions, forces cancel and R vanishes, whereas any finite value signals that wake-mediated asymmetry injects net momentum into the system. This measure has the advantage of being experimentally accessible from trajectory data, avoiding the need for coarse-grained approximations or indirect stability proxies.
@@
Figure 9: Non-reciprocal driving strength $R$ versus bias voltage. Shaded regions show the distribution of $R$ values, while points and error bars denote the spread of the data.
@@
We have calculated the value of R for all particle pairs, and the results show a clear dependence on the confinement bias voltage. Fig. 9 illustrates how the distribution of R varies with changes in confinement voltage. In the figure, the shaded regions represent the spread of pairwise non-reciprocal driving strengths at each voltage. A wider shaded region indicates greater variation in R among particle pairs. As the confinement voltage decreases, the spread of the distribution increases, indicating enhanced non-reciprocity and variability. At 140 V, the measured R is small, indicating that interlayer forces are relatively reciprocal and the crystal retains long-range order. Reducing the confinement to 120 V increases interlayer overlap and enhances wakefield-mediated attraction, raising the value of R. This coincides with a marked increase in particle-pairing events. At 110 V, where the bilayer collapses into a strongly fluctuating state, the nonreciprocal driving strength reaches its maximum, confirming that asymmetric forces dominate the dynamics at low confinement. The monotonic growth of R across this range establishes a direct experimental link between microscopic nonreciprocity and macroscopic melting.
@@
The importance of this observation is underscored by theoretical work by Ivlev et al.[42], which demonstrates that nonreciprocal interactions fundamentally alter statistical mechanics. Unlike equilibrium Yukawa crystals, where equipartition ensures a well-defined temperature, nonreciprocal systems can sustain different effective kinetic temperatures for different particle species, or even exhibit continuous energy growth with no steady-state equilibrium. Our data support the latter scenario: as non-reciprocity increases, particle pairs act as channels for energy injection from the ion flow into the lattice. The repeated formation and collapse of pairs provides the necessary energy to local rearrangements and velocity fluctuations, ultimately destabilizing the entire crystal.
@@
An additional consequence of non-reciprocity is the emergence of a dynamical asymmetry in how instabilities develop. Classical Mode-Coupling Instability (MCI) is triggered when longitudinal and vertical phonon modes intersect, producing resonant energy transfer. In contrast, in a bilayer system, the phonon spectra exhibit MCI-like features (frequency shifts, harmonics) but without the expected resonance crossing of longitudinal and vertical modes. This indicates that the instability originates from the combined action of wake-driven non-reciprocity and modified mode coupling, rather than classical MCI alone.
@@
In summary, the analysis of nonreciprocal interactions demonstrates that melting in bilayer dusty plasmas is not only driven by damping thresholds or classical instabilities, but also by a confinement-controlled increase in wake-mediated asymmetry. The monotonic rise of nonreciprocal driving strength with decreasing confinement voltage provides a quantitative, experimentally validated mechanism for how crystalline order gives way to fluid-like coexistence. This establishes nonreciprocity not only as a descriptive feature of bilayer plasmas but as the primary physical agent responsible for their phase transition.

### v3

@@ Abstract
This article presents a detailed investigation of solid-fluid phase coexistence in a bilayer dusty plasma crystal subjected to varying confinement ring bias voltages in a DC glow discharge argon plasma. Melamine formaldehyde particles were employed to form a stable, hexagonally ordered bilayer crystal within a confinement ring electrically isolated from the grounded cathode. By systematically adjusting the confinement ring bias, a distinct phase coexistence emerged characterized by a fluid-like melted core surrounded by a solid crystalline periphery. Crucially, analysis of the phonon spectra revealed frequency shifts that deviate significantly from the predictions of classical monolayer Mode-Coupling Instability (MCI) theory. Stability analysis further demonstrated that dynamic interlayer particle pairing and the associated increase in non-reciprocal interaction strength are strongly correlated with the onset of structural destabilization. These findings highlight previously underappreciated mechanisms driving the melting transition in bilayer dusty plasmas, offering a more comprehensive understanding of phase behavior in complex plasma systems. The results underscore the importance of interlayer coupling and confinement effects in tuning structural transitions.
@@ I Introduction
In recent years, dusty plasma crystals have emerged as a model system for exploring phase coexistence [15, 16, 17, 18, 19, 20, 21], where solid and fluid phases can coexist within the same structure. The dynamics underlying this phenomenon are complex, as they involve particle interactions, structural rearrangements, and instabilities that are not commonly observed in simpler systems. Phonon modes play a critical role in understanding phase transitions and melting dynamics in complex plasmas. Longitudinal and transverse phonon modes, arising from compressional and shear motions respectively, provide insights into system stability and energy distribution [22, 23, 24]. Wang et al. [22] developed a theoretical framework for phonon modes in Yukawa crystals, a model analogous to dusty plasmas with screened Coulomb interactions. This enabled systematic studies of dispersion relations and the influence of confinement and interaction strength on phonon behavior and instabilities  [25, 26, 7, 27].
@@
Studies on phase transitions in monolayer complex plasmas have shown that mode-coupling instability (MCI) is the primary mechanism responsible for melting.  [35, 7, 26, 34, 27]. It was observed that the coupling between longitudinal and out-of-plane modes triggers the destabilization of the lattice, ultimately leading to melting. Ivlev et al. [36] later proposed a common framework for monolayer and bilayer crystals, highlighting wake-mediated non-reciprocity, MCI and a universal stability principle for both monolayer and bilayer complex plasmas. The instability is driven by the coupling of wave modes, which is caused by non-reciprocal, wake-mediated interparticle forces. In monolayers, ion wakes introduce a weak negative coupling between the normally separate horizontal and vertical modes. Conversely, bilayers possess an inherent positive coupling between horizontal modes due to their structural asymmetry. Therefore, for a bilayer to become unstable, the wake-induced negative coupling must be strong enough to overcome this existing positive offset, a condition that is facilitated in both systems when modes are in resonance. Recently, Jaiswal et al. [16] reported the first experimental observation of phase coexistence in bilayer dust clouds, induced by confinement variations. While they observed signs of MCI, the absence of intersection between vertical and horizontal modes and phase coexistence seemed to emerge when interlayer and interparticle spacing approach or fall below the screening length. This highlights the need for detailed investigation of wake-mediated non-reciprocal interactions and their role in bilayer crystal stability.
This article investigates the mechanisms underlying phase coexistence and melting in bilayer dusty plasma crystals, with a specific focus on the interplay between particle pairing and mode coupling under varying confinement ring bias voltages in a DC glow discharge plasma. By analyzing phonon spectra across a range of confinement potentials, we explore how longitudinal and transverse modes respond to particle interactions and coupling effects. Through detailed tracking of particle trajectories, separations, and velocities, we uncover new insights into the dynamic processes that destabilize the crystal structure. Our results demonstrate that dynamic particle pairing, combined with a modified form of MCI, is strongly correlated with the observed structural transition of bilayer dust crystal. Additionally, we introduce and apply a quantitative analysis of interlayer non-reciprocity, capturing the emergence of asymmetric particle interactions that signal structural instability. This approach provides a novel diagnostic for identifying the onset of disorder and melting in bilayer dusty plasmas. Overall, this study contributes to a deeper understanding of structural stability and phase transitions in complex plasma systems, offering new perspectives on the microscopic interactions that govern these phenomena.
@@ III Results
Fig. 2 shows an overlapped image of 20 consecutive frames from the top view camera at 103 V of confining ring voltage. The circular marked region shows a clear melting whereas the periphery is almost stationary with a small thermal fluctuation around the equilibrium position. The side view of the particle cloud (shown in Fig. 3) reveals that the interlayer gap between two layers is decreasing as the confinement potential decreases from 140 V to 103 V. This increases the interaction between the layers and vertical oscillations and the coupling of the two layers was seen for the voltage below 120 V (see supplementary material).
Clear vertical oscillations in both layers suggest that the Schweigert mechanism may not be directly applicable to our bilayer system. Moreover, this mechanism assumes a damping-strength threshold beyond which the crystal melts, whereas in our case the damping strength (i.e., neutral pressure) remains constant. Our results indicate that structural changes are mainly driven by variations in the confining potential, so the damping threshold appears to play little role in this context. Side-view imaging provides further insight into the structural evolution of the bilayer with decreasing confinement bias. At high bias voltages (140 V), the system forms a stable bilayer with small-amplitude vertical oscillations in both layers. As the bias voltage is reduced toward 120 V, vertical oscillations become more pronounced in both layers (see Supplementary Figs. 2 and 3), indicating increased coupling to the sheath electric field. Upon further reduction of the bias voltage, the upper layer intermittently splits into two vertically separated sublayers, while the lower layer continues to oscillate as a single layer. The splitting of the upper layer causes the reduced layer gap and enhances both interlayer and intralayer interactions. The combined effect of strengthened vertical oscillations in both layers and the fragmentation of the upper layer promotes particle pairing and non-reciprocal energy transfer, ultimately leading to the melting observed in the top-view measurements.
@@ III.1 Phonon spectra
The phonon spectrum of the dust lattice [35, 33] was analyzed over a range of confinement ring bias voltages, focusing on optical modes (from side-view data) and acoustic modes (from top-view data). Particle positions were extracted using ImageJ [37] for both views.
The extracted positions were processed using the Trackpy library [38, 39] to link particle trajectories and compute velocities. Longitudinal and transverse phonon currents were then calculated following the method of Nunomura et al. [23], and Fourier transforms were applied to obtain the current fluctuation spectra. This analysis across various ring bias voltages enabled us to study the evolution of wave modes in the crystal.
@@
At a confinement bias of 120 V, signs of torsional motion [40] become visible, indicating the onset of particle pairing [see hotspots in Figs. 5(c), (g) and 5(c), (g)]. As previously reported, torsion and MCI develop over similar spatial and temporal scales [40]. However, in contrast to monolayer MCI, no localized hot spots are present in the out-of-plane spectra (Fig. 6(c)). Below 120 V, spectral features appear increasingly blurred. It is possible that the imaging setup captured particles from both layers, making the linking of particle trajectories inaccurate and leading to artificial mixing of spectral features. This behavior likely reflects a change in interparticle coupling, possibly due to the formation of particle pairs between the two layers.
Such pairing can modify the system’s vibrational response and affecting the mechanical stability of the lattice [40].
@@ III.2 Dependence of particle pairing on the confining ring bias voltage
As the ring bias voltage was reduced, a noticeable increase in the number of particle pairs formed was observed. Tracking the number of pairs as a function of voltage revealed a clear trend: the number of pairs increased as the voltage decreased (Fig. 7), particularly around 120 V and below. At 103 V, the number of pairs in the field of view was measured as 107 ± 10, confirming the strong increase in pair formation at lower confinement potentials
@@ III.3 Case study: dynamics of particle pairing at a confining ring bias voltage of 120 V
As we have seen in previous sections, at 120 V, significant changes in particle behavior were observed. For example, the hot spot arises at 120 V and the pairing effect between the particles also increases drastically after 120 V. This has important consequence for particle structural dynamics therefore, 120 V has been considered as a critical voltage for understanding the transition from crystalline to fluid states.
In order to illustrate that, we focus on three particles: "1" and "2" in the top layer (Layer 1 in Fig. 8(a)), and "3" just above Layer 1. Their motion over 6 seconds reveals the dynamics of pair formation, collapse, and particle dragging, contributing to system destabilization. Fig. 8(b) shows the evolution of their $x$-coordinates.
@@
Initially, particle "3" formed a stable pair with "1" for $\sim$1.5 s, as shown by their nearly identical $x$-coordinates. At $t\simeq 1.5$ s, the pair broke, and "3" drifted leftward toward particle "2," forming a new pair shortly before $t=2$ s. Around $t\simeq 3$ s, particle "3" continued leftward and while dragging along particle "2" which gradually moved leftwards too. The pair realigned by $t\simeq 4$ s. This dragging is the signature of the ion wake below particle "3" influences nearby particles (also evidenced by Melzer et al.[30]), potentially destabilizing the entire system.
@@
where $\gamma$ is the Epstein drag coefficient [41], $\langle V\rangle$ is the average horizontal velocity and $m$ is the mass of the particles.
Assuming the system is overdamped and using $F_{x}=k\cdot\Delta x$, the spring constant for the 2-3 pair during the interaction was:
@@
In this expression, $\Delta x$ is defined as the instantaneous horizontal displacement of the particle from its equilibrium position, corresponding to the effective lateral deformation produced by the wake-induced interparticle attraction. The estimated horizontal dragging force and effective spring constant quantify the strength of the wake-mediated coupling between paired particles. These values confirm that the interparticle attraction is sufficiently strong to produce measurable lateral displacement and energy exchange. The spring constant thus characterizes how mechanical energy is stored and released during pairing events, linking microscopic force asymmetry to the observed velocity fluctuations and, ultimately, to the onset of melting. [34].
@@
III.4 non-reciprocal interactions and driving strength
@@
To quantify this effect, we introduced a pair-resolved metric following from the recent work from Ivlev et al. [42] {align} R=<|F_top+F_bottom|>
@@
Here $F_{top}$ and $F_{bottom}$ denote the instantaneous forces acting on the upper and lower layer particles within a pair, respectively, as obtained from their measured accelerations using the particle trajectories. This directly probes the imbalance of forces in paired particles across the two layers. In the case of reciprocal interactions, forces cancel and R vanishes, whereas any finite value signals that wake-mediated asymmetry injects net momentum into the system. This measure has the advantage of being experimentally accessible from trajectory data, avoiding the need for coarse-grained approximations or indirect stability proxies.
@@
Figure 9: Non-reciprocal driving strength $R$ versus bias voltage in the region where melting occurs. Shaded regions show the distribution of $R$ values, while points and error bars denote the spread of the data.
Figure 10: Non-reciprocal driving strength $R$ versus bias voltage at the periphery. Shaded regions show the distribution of $R$ values, while points and error bars denote the spread of the data.
@@
We have calculated the value of R for all particle pairs, and the results show a clear dependence on the confinement bias voltage. Fig. 10 and Fig. 10 illustrate how the distribution of R varies with changes in confinement voltage in the central region that is melting and the periphery respectively. In the figures, the shaded regions represent the spread of pairwise non-reciprocal driving strengths at each voltage. A wider shaded region indicates greater variation in R among particle pairs. As the confinement voltage decreases, the spread of the distribution increases for melted region (Fig 10), indicating enhanced non-reciprocity and variability while the driving strength remains relatively unchanged from the periphery (Fig. 10).
At 140 V, the measured R is small, indicating that interlayer forces are relatively reciprocal and the crystal retains long-range order. Reducing the confinement to 120 V increases interlayer overlap and enhances wakefield-mediated attraction, raising the value of R. This coincides with a marked increase in particle-pairing events. At 110 V, where the bilayer collapses into a strongly fluctuating state, the non-reciprocal driving strength reaches its maximum, confirming that asymmetric forces dominate the dynamics at low confinement.
The monotonic growth of R accompanies the onset of melting and indicates that strengthening non-reciprocal interactions likely contributes to destabilization of the lattice, in combination with confinement-driven modifications of local interaction forces.
It is worth mentioning that the neutral gas pressure and discharge voltage were kept constant throughout the experiment; therefore the global plasma parameters remained effectively unchanged. The ring bias modifies only the local sheath electric field in the vicinity of the dust cloud, producing local changes in ion streaming speed, Debye screening and wake-mediated asymmetry. Consequently, the observed changes in R reflect local interaction modifications rather than global plasma reconfiguration.
@@
The importance of this observation is underscored by theoretical work by Ivlev et al.[42], which demonstrates that non-reciprocal interactions fundamentally alter statistical mechanics. Unlike equilibrium Yukawa crystals, where equipartition ensures a well-defined temperature, non-reciprocal systems can sustain different effective kinetic temperatures for different particle species, or even exhibit continuous energy growth with no steady-state equilibrium. Our data support the latter scenario: as non-reciprocity increases, particle pairs act as channels for energy injection from the ion flow into the lattice. The repeated formation and collapse of pairs provides the necessary energy to local rearrangements and velocity fluctuations, ultimately destabilizing the entire crystal.
@@
An additional consequence of non-reciprocity is the emergence of a dynamical asymmetry in how instabilities develop. Classical Mode-Coupling Instability (MCI) is triggered when longitudinal and vertical phonon modes intersect, producing resonant energy transfer.
In contrast, in a bilayer system, the phonon spectra exhibit MCI-like features (frequency shifts, harmonics) but without the expected resonance crossing of longitudinal and vertical modes. This indicates that the instability originates from the combined action of wake-driven non-reciprocity and modified mode coupling, rather than classical MCI alone.
@@
In summary, the analysis of non-reciprocal interactions demonstrates that melting in bilayer dusty plasmas is not only driven by damping thresholds or classical instabilities, but also by a confinement-controlled increase in wake-mediated asymmetry. The monotonic rise of non-reciprocal driving strength with decreasing confinement voltage provides a quantitative, experimentally validated mechanism for how crystalline order gives way to fluid-like coexistence.
These findings show that increasing non-reciprocal interaction strength is strongly associated with the onset of melting and likely plays a significant role in destabilization, together with other confinement-induced changes in local sheath structure and particle coupling.